OpenStudy (anonymous):

3.Solve x over 2 + 4 = 10.

OpenStudy (anonymous):

do you mean (x/2)+4=10 if so it would be (x/2)+(4/1)=10 (4/1)(2/2)=8/2 so the denominators are equal (x/2)+(8/2)=10 (x+8/2)=10 x+8=20 x=12